# Visualizing Causality in Mixed Reality for Manual Task Learning: A Study

### v3

@@ I Introduction
Derived from these prior works, research pointed out that learners can anticipate their future steps [15, 16], connected to their current steps. Similarly, discussion in the psychology community holds that human learns tasks by understanding the causality, i.e., the cause and effect relations among the steps that lie within the tasks [17, 18, 19, 20, 21]. Enlightened by these findings, prior works have also explored the visualization of causality in various methods such as pre-cueing [22, 15, 23, 24] or visualizing past or future events [25, 26] in MR. For example, Liu et al. [15] suggest pre-cueing for manual tasks, providing cues before the actual event occurs, which can prepare the user for the upcoming task. Similarly, showing past and future events in the task can help humans understand the causality within the task [25]. It is also manifested how understanding a task’s cause and effect helps prevent errors in future steps [22].
Prior works have focused on diverse applications of visualizing causality in MR and, therefore, enlightened our discussion over the effect of visualization of causality in MR on the human learning process, particularly in our scope, the manual task learning process. Based on our literature reviews, we categorize the MR visualization of causality in manual tasks into three levels: gesture, interaction, and event. Gesture-level causal visualization reveals the details of hand-object interaction that are required for future steps, e.g. angles, amount of translation, directions, etc. An interaction defines a set of dynamic gestures with an object, and interaction-level causal visualization displays the causal relationships between the interactions. An event defines a set of interactions with different objects, and event-level causal visualization reveals the causal relationships among the events. Although different levels of causal visualization have been implemented among the prior works, there exists no systematic study to reveal the effect of visualizing causality in MR on the learning gain of manual tasks. Moreover, there exists no discussion over the potential differences among different methodologies of causal visualization in MR.
@@
To answer these questions, we built a three-level hierarchy for causality representation in manual tasks, deriving from the existing psychology research on causality and human intention [27, 28], namely event-level, interaction-level, and gesture-level, corresponding to three levels of human-intention and causality. Based on the derived hierarchy, our embodied demonstrations followed four causal visualization options: no causality, task-level causality, interaction-level causality, and gesture-level causality, where the users are shown the corresponding demonstrations of manual tasks that are causally connected to their current step.
To investigate the effects of these four options on the learning gain in manual task learning in MR, we conducted two sessions of studies ($N=48$), where we embodied the step-by-step demonstration of the hand-object avatars in a manual assembly task in users’ MR view while they were learning the task with the physical components in situ. The contributions of our papers are as follows:
@@
  • Quantitative and Qualitative Results showing users’ objective/subjective responses and learning gains from the sessions of manual tasks while following different visualizations of causality.
  • Design Recommendation and In-depth Discussion summarized from the results and further discussion of the study, suggesting promising directions that deepen the research in manual task learning in MR.
@@ III-A Hierarchy of Causality in Task
Moreover, intentions not only decide the occurrences but also decide the order and pattern of the elements in the inter-layer(i.e., they are only connected in the same layer). These interlayer connections between elements are related by cause and effect(causality) [94]. This implies that the intentions themselves drive the causality observed at each level of the hierarchy, and this intention-driven causality Figure 1. includes 1) Event level causality 2) Interaction level causality and 3) gesture level causality.
@@ V User Study
We conducted a two-session user study with our test bed. In the first session, we studied the impact of the existence of causality visualization on the learning process of a manual task. In the second session, we studied the impact of different levels of causality visualization on the learning process.
@@ V-A Hypotheses
Based on our research questions and literature review, we formulated the following hypotheses for Session 1:
@@ V-B1 Participants
We evaluated the learning performance by modeling the causality for Groups (G1, G2, and G3), while the Control group (GC) was provided no information on causality. This approach enabled comparison of the learning performance between participants exposed to different levels of our taxonomy for modeling causality. In our study, we would like to clarify that the understanding of causality has a cliff-like learning effect, i.e., ”know if a thing will happen or not,” is an approximate binary variable. We design our user study with a between-subject setup.
@@ V-B2 Procedure
Each participant was warmly welcomed by the researcher and provided with a concise introduction to both the study and the task. Upon signing the consent form, participants received a brief overview of the task’s details. Then, they were asked to wear headsets with pass-through capabilities to get familiar with our system in the MR environment. The study includes two phases: the learning phase and the testing phase. In the learning phase, each user was provided with sufficient time (1 hour) to learn the content using our system. Every participant finished learning within one hour. Then, participants performed the entire assembly task without any help in the testing phase. This phase has a 15-minute time limit to complete the task. After finishing the study, users provided feedback about the study through a semi-structured interview and questionnaires.
@@ V-B3 Data Collection
Throughout the study, we recorded participants’ performances from the third-person camera as well as a screen capture of the AR view during the session. We then manually computed the length of time of each session, event, and interaction alongside task/sub-task accuracy in post-processing. We visually compared the procedural consistency between participants’ demonstrations of tutorials at all three levels. We analyzed the number of times each participant from Gc and G3 revisited the instruction, i.e., they referred to the instruction again for more information. We did not include participants from G1 and G2 since the coverage in the length of the instructions is different. Therefore, it is of no need or rigor to compare with them.
After the study, participants provided feedback on our system through the NASA TLX [98] form and a semi-structured interview, Participants also completed a 5-point Likert scale questionnaire. The questions from the survey were designed to determine the participant’s memory of the task, self-rating of their memory from learning, motivation to learn, and comprehension of the causality and intention within the tasks.
@@ VI Results and Data Analysis
Thus, we make four comparisons (Gc vs G3, G1 vs G2, G1 vs G3, and G2 vs G3). It is shown by the Shapiro-Wilk normality test ($p<0.005$) that the data from neither our quantitative nor qualitative evaluation follows a normal distribution. Therefore, we conducted a Friedman test for G1 vs G2 vs G3 and a Wilcoxon signed-rank test for all four comparisons individually. To prove the hypothesis of session 1, we compare (Gc vs G3 and for session 2 the other three comparisons.
@@ VI-A1 Learning Performance
The Friedman test shows significant differences among the groups in learning time($\chi^{2}(3)=18.725,p=0.003$). The average learning time shows that the participants spend the longest amount of time (all in minutes) when provided event-level causal demonstration in G1($M=24.71,SD=3.08$). To evaluate H1, we compare the Gc and G3.The average learning time of participants from G3 ($M=22.27,SD=3.23$) spend more time than those from Gc ($M=18.24,SD=2.60,p=0.001,Z=-2.981$). This shows that G3 which includes causal information increases the learning (training) time as compared to Gc which shows only the current task. Hence H1 is failed. For H4, we compare among G1, G2 and G3. Participants from G2 ($M=21.26,SD=2.71$) spend insignificantly more time ($p=0.140,Z=-1.098$) than those from G3. Meanwhile, participants from G2 spend significantly less time in learning than those from G1 ($p=0.008,Z=-2.401$). Participants from G1 also spent significantly more time learning than those from G3 ($p=0.050,Z=-1.647$). We conclude that H4 is partly proved as the most time was taken by G1 and least by G2
@@ VI-A2 Testing Performance
We evaluate the remaining four hypotheses H2, H3, H5, and H6 in this section. The participant’s performance during the two testing phases with respect to the average total task time of the first phase is shown in  Figure 9 (H2 and H5). The number of mistakes of three different levels made during both phases is demonstrated in  Figure 10 (H3 and H6).
The Friedman test shows significant differences among the groups in testing time ($\chi^{2}(3)=15.9,p=0.001$) The average testing time of the participants from G3($M=4.31,SD=0.35$) is significantly lower than all three other groups, namely Gc($M=5.20,SD=0.69,p=0.002,Z=-2.824$), G1 ($M=4.92,SD=0.47,p=0.001,Z=-2.981$), and G2 ($M=4.61,SD=0.24,p=0.036,Z=-1.804$). This proves H2 and H5. The average testing time of the participants from G1 is insignificantly lower than those from G2 ($p=0.058,Z=-1.5689$).
@@ VI-B Subjective Feedback
This also proves H5 We also evaluated the number of mistakes made by the participants during the testing phase. The Friedman tests show no significant difference among the groups in total mistake numbers($\chi^{2}(3)=3.3,p=0.348$), event-level mistake numbers ($\chi^{2}(3)=0.5,p=0.919$), event-level mistake numbers ($\chi^{2}(3)=5.825,p=0.120$), and event-level mistake numbers ($\chi^{2}(3)=4.525,p=0.210$). This shows that both H3 and H6 fails.
@@ VI-B1 Likert-Scale Questionnare
  • Understanding. The Friedman test shows a significant difference between all four groups for Q1($\chi^{2}(3)=8.525,p=0.036$) and Q2($\chi^{2}(3)=8.05,p=0.044$) but not Q3($\chi^{2}(3)=0.175,p=0.98$). We performed pairwise comparisons among the groups for Q1 and Q2. G3 ($M=4.42,SD=0.75$) understands the importance of the task better as compared to G1 ($M=3.92,SD=1.21,p=0.020,Z=-2.038$) and Gc ($M=2.58,SD=1.24,p=0.008,Z=-2.400$). G3 ($M=3.92,SD=0.9$) also knows better the outcome of their actions beforehand as compared to Gc ($M=2.75,SD=1.35,p=0.038,Z=-1.765$)  Figure 11.
@@ VI-B2 NASA TLX
We used NASA TLX’s responses to evaluate the perceived subjective mental workload experience for the sessions. The findings are present in  Figure 14. After confirming the non-normality with a Shapiro-Wilk test, a Friedman test was conducted on the results obtained from an unweighted NASA TLX to compare all the sessions. None of the metrics showed significance: Mental Demand ($\chi^{2}(3)=3.875,p=0.275$), Physical Demand ($\chi^{2}(3)=1.8,p=0.614$), Temporal Demand($\chi^{2}(3)=4.225,p=0.238$), performance ($\chi^{2}(3)=2.625,p=0.453$), Effort($\chi^{2}(3)=3.525,p=0.317$) and frustrations($\chi^{2}(3)=2.125,p=0.546$). We further do not perform any adjustments in p-values for multiple comparisons, as there is no significant result as shown in  Figure 14.
@@ VI-C Research Question Analysis
  • RQ1: We observed that H1 and H3 did not yield successful outcomes, whereas H2 was confirmed to be accurate. H1 posited that the learning time in G3 would surpass that of Gc. After failing of two hypotheses we further analyzed more results. To explore performance improvements during the learning phase, we incorporated an additional metric based on the number of tutorial revisits [6]. Interestingly, the average revisits for Gc ($M=3.17,SD=0.72$) were significantly higher compared to G3 ($M=2.33,SD=0.49,p=0.018,Z=-2.090$), as depicted in  Figure 8. This shows the difficulty in learning the content. Regarding H3, it was inconclusive due to the absence of a significant difference. We conclude that the absence of significance in the number of mistakes is due to the similar mistakes made during testing, which led to close counts of mistakes across the groups at all three levels.
@@
  • RQ2: We can see from the results that H4 is rejected although G1 did take the most time which corresponds to event-level causality. H5 was proven valid whereas H6 is not supported. From further analyzing the data we found that the reason behind the failure of H6 is the same as discussed for H3 in the above paragraph for the mistakes made in the task. Overall, we can conclude that the complete hierarchy of causality with all three levels improves the performance of the task.
@@ VII-A Better Learning Results with Intention-driven Causality
VII-B Active Inference of Causality
Although we have discovered the gap of understanding in intention-driven causality of their actions between Gc and G3 ( Understanding - Q1,Understanding - Q2), it is interestingly worth noticing that both groups show insignificant difference in their comprehension of the overall goal (Understanding - Q3). As suggested in  [89, 90, 91], learners can infer the causality and the intention of the demonstrator while observing. We anticipate that during our study, learners will be also able to actively infer the causality within the task, in addition to passively learning causality from the causal information provided by the system. Between G2 and G1, there exist significant gaps in the understanding of causal relation, indicating that participants from G1 (who were not provided with interaction level causality information) cannot infer interaction level causality as much as that we provide to the participants from G2. We summarize that learners are not able to fully acknowledge the causality connecting their current interaction, although they can infer the event-level causality by observational learning. Moreover, it remains an interesting open research question to what level and how much of the causality the learners can infer by observation. We give details of future work in this direction in  subsubsection VIII-A3.
VII-C Levels of Details: the Deeper, the Better
@@
From our cross-comparisons among G1, G2, and G3, we notice that both G2 and G3 have significantly less learning time and less testing time than G1. The higher learning time of G1 comes from the fact that showing only event-level instruction makes it hard for the learners to memorize the procedure and results in the learners more frequently referring to the instructions. ”The instruction is too long for me to remember, and I have to rewind to the part I am doing (G1, P3).” Meanwhile, both G2 and G3 have less numbers of mistakes during their demonstrations. The ratings from the question Q1 show that G1 do not possess as much causal understanding as G3, which indicates that showing only the event level intention-driven causality does not provide a deeper knowledge of the causal relations within the tasks. The feedback from Memory - Q2 also indicates that without learning the interaction and gesture/affordance levels of intention-driven causality, G1 has a poorer understanding of how to operate the tool. A similar conclusion can be drawn from Memory - Q3 where G1 gives lower ratings on their performance than G2 and G3 regarding their precision. Although the Friedman test revealed insignificant differences among the group in responses to Memory - Q3, we conducted pair-wise Wilcoxon signed rank tests for G2 vs G1 ($p=0.018,Z=-2.073$) and G3 vs G1 ($p=0.023,Z=-1.988$) just to validate the similar conclusion from Memory - Q2.
Participants from G2 spend insignificantly less time in the learning phase than those from G3, since G2 are not learning the additional information about the gesture/affordance level causality. However, G3 have a significant reduction in testing time and fewer total mistakes during testing.
@@
The results above indicate that learners’ efficiency and understanding are positively correlated with more detailed intention-driven causality information, although compromises are made in terms of the insignificant amount of extra time spent on learning. Prior works [101, 8, 102] have shown that the more details provided to the users of the MR system, the higher cognitive load is forced on the users, resulting in decreasing learning performance. Yet, with details added in our study, we do not observe a significant increase in learning time or a significant difference in users’ subjective confidence about the task they have learned (Self-confidence - Q1,2,3). To sum up, we conclude that including causal information in manual task learning in MR provides the learners with more and better learning gains, however, might not be the best practice when efficiency in learning is the priority.
VII-D Neglecting Details
An interesting finding from the numbers is that all four groups give similarly high ratings in Understanding - Q3, indicating that regardless of the methodology applied, all learners show a good understanding of the overall goal of the task. Likewise, all groups report high ratings of their confidence in performing the task in Self-confidence - Q1. Moreover, G2 and G3 show similar confidence in their tool operation (Self-confidence - Q2) and task precision (Memory - Q3). Yet their performance in time and the number of mistakes made at the gesture level are significantly different (Despite the insignificance among the groups shown by the Friedman test, we ran an individual comparison test for gesture-level mistakes ($p=0.034,Z=-2.0226$)). This contradiction can be best explained by a lack of understanding. Learners exposed to more detailed causality and intent information had a deeper understanding of the process, enabling a more complete picture of how tools and actions contribute to the goal. By seeing the future that is causally connected to their present, they can obtain more knowledge about their present. With this knowledge, comes the ability to ascertain (1) which parts of the details of the instructions are causally linked, (2) which parts of the details can be safely ignored, and (3) which parts of the details are most important for efficient, accurate task completion ”I saw that I was about to attach a camera, then I just guessed that I needed to make the hole big enough for the bolts to go through (G3, P9).
However, we further argue that even in G3, where all intention-driven causality information is presented, learners still neglect important details due to the static nature of our implementation. In other words, the causality is presented to the learners by merely showing the ideal future state instead of showing how the future is being affected by the present. ”I knew I would install the arm later onto the stand, but I didn’t know I had put the stand backward (which was wrong) until I was about to install the arm (G3, P1).” To address this problem, in  subsubsection VIII-A1 we propose a method for learners to immediately understand their errors by predicting the action’s influence on the intention.

### v4

@@ I Introduction
Derived from these prior works, research pointed out that learners can anticipate their future steps [15, 16] , connected to their current steps. Similarly, discussion in the psychology community holds that human learns tasks by understanding the causality, i.e., the cause and effect relations among the steps that lie within the tasks [17, 18, 19, 20, 21]. Enlightened by these findings, researchers have also explored the visualization of causality in various methods such as pre-cueing [22, 15, 23, 24] or visualizing past or future events [25, 26] in MR.
Prior works have focused on diverse applications of visualizing causality in MR and, therefore, enlightened our discussion over the effect of visualization of causality in MR on the human learning process, particularly in our scope, the manual task learning process. Based on our literature reviews, we categorize the MR visualization of causality in manual tasks into three levels: gesture, interaction, and event. While prior research has explored the use of MR for manual task learning and the visualization of causality to some extent, no systematic study has examined how different levels of causal visualization: gesture level, interaction level, and event level impacts learning outcomes. Specifically, there has been no comparative analysis of these methods, leaving a critical gap in understanding how to optimize causal visualization for effective manual task learning in MR environments.
@@
To answer these questions, we built a three-level hierarchy for causality representation in manual tasks, deriving from the existing psychology research on causality and human intention [27, 28], namely event-level, interaction-level, and gesture-level, corresponding to three levels of human-intention and causality. Based on the derived hierarchy, our embodied demonstrations followed four causal visualization options: no causality, event-level causality, interaction-level causality, and gesture-level causality, where the users are shown the corresponding demonstrations of manual tasks that are causally connected to their current step.
To investigate the effects of these four options on the learning gain in manual task learning in MR, we conducted a two-phase study ($N=48$), where we embodied the step-by-step demonstration of the hand-object avatars in a manual assembly task in users’ MR view while they were learning the task with the physical components in situ. The contributions of our papers are as follows:
@@
  • Quantitative and Qualitative Results showing users’ objective/subjective responses and learning gains of the manual tasks while following different visualizations of causality.
  • Design Recommendation and In-depth Discussion summarized from the results and future opportunity suggesting promising directions that deepen the research in manual task learning in MR.
@@ III-A Hierarchy of Causality in Task
Moreover, intentions not only decide the occurrences but also decide the order and pattern of the elements in the inter-layer(i.e., they are only connected in the same layer). These inter-layer connections between elements are related by cause and effect(causality) [94]. This implies that the intentions themselves drive the causality observed at each level of the hierarchy, and this intention-driven causality Figure 1. includes 1) Event level causality 2) Interaction level causality and 3) gesture level causality.
@@
For example, consider the task of assembling furniture. At the event level, the task could involve ”assembling the frame,” ”attaching the shelves,” and ”tightening the screws.” These events must occur in a specific sequence, the frame must be assembled before attaching the shelves. The causality here is driven by the assembler’s P-Intention to complete the furniture assembly.
At the interaction level, within the event of ”assembling the frame,” interactions could include ”positioning the side panels,” ”aligning the bolts,” and ”inserting the bolts into the frame.” These interactions are causally linked, positioning the side panels must occur before aligning and inserting the bolts. The assembler’s D-Intention to successfully complete the frame event governs this level of causality.
At the gesture level, during the interaction of ”inserting the bolts,” gestures such as ”grasping the bolt,” ”aligning the bolt with the hole,” and ”turning it with a wrench” come into play. These gestures are causally connected, where misalignment in one gesture can disrupt the interaction. These gestures are guided by the M-Intention to complete the interaction correctly.
@@ V User Study
We conducted a user study with our test bed to evaluate two evaluate our research questions. In the first, we studied the impact of the existence of causality visualization on the learning process of a manual task. In the second, we studied the impact of different levels of causality visualization on the learning process.
@@ V-A Hypotheses
Based on our research questions and literature review, we formulated following hypothesis based on time and errors which are well-established factors in assessing performance in manual task learning.
For RQ 1:
@@ V-B1 Participants
We evaluated the learning performance by modeling the causality for Groups (G1, G2, and G3), while the Control group (GC) was provided no information on causality. This approach enabled comparison of the learning performance between participants exposed to different levels of our taxonomy for modeling causality. Testing causality levels hierarchically mirrors how learners typically develop their understanding level by level. This approach allows us to evaluate how incremental layers of causality improve task learning, rather than isolating them as independent conditions. Therefore, we chose to test causality levels hierarchically instead of independently. In our study, we would like to clarify that the understanding of causality has a cliff-like learning effect, i.e., ”know if a thing will happen or not,” is an approximate binary variable. We design our user study with a between-subject setup.
@@ V-B2 Procedure
Each participant was warmly welcomed by the researcher and provided with a concise introduction to both the study and the task. The study was conducted under the approval of the Institutional Review Board. Upon signing the consent form, participants received a brief overview of the task’s details. Then, they were asked to wear headsets with pass-through capabilities to get familiar with our system in the MR environment. The study includes two phases: the learning phase and the testing phase. In the learning phase, each user was provided with sufficient time (1 hour) to learn the content using our system. Every participant finished learning within one hour. Then, participants performed the entire assembly task without any help in the testing phase. This phase has a 15-minute time limit to complete the task. The pre-authored content is visualized in the UI panel in the form of VR scene snapshots as exemplified in  Figure 6. Participants can also see AR demonstrations of the current interaction as exemplified in Figure 7. After finishing the study, participants provided feedback about the study through a semi-structured interview and questionnaires.
@@ V-B3 Data Collection
Throughout the study, we recorded participants’ performances from the third-person camera as well as a screen capture of the AR view. We then manually computed the length of time of each phase, event, and interaction alongside task/sub-task accuracy in post-processing. We visually compared the procedural consistency between participants’ demonstrations of tutorials at all three levels. We analyzed the number of times each participant from Gc and G3 revisited the instruction, i.e., they referred to the instruction again for more information. We did not include participants from G1 and G2 since the coverage in the length of the instructions is different. Therefore, it is of no need or rigor to compare with them. After the study, participants provided feedback on our system through the NASA TLX [98] form and a semi-structured interview, Participants also completed a 5-point Likert scale questionnaire. The questions from the survey were designed to determine the participant’s memory of the task, self-rating of their memory from learning, motivation to learn, and comprehension of the causality and intention within the tasks [99, 100], we designed the questionnaires following prior works on learning manual tasks in MR [8, 36, 3].
@@ VI Results and Data Analysis
Thus, we make four comparisons (Gc vs G3, G1 vs G2, G1 vs G3, and G2 vs G3). It is shown by the Shapiro-Wilk normality test ($p<0.005$) that the data from neither our quantitative nor qualitative evaluation follows a normal distribution. Therefore, we conducted a Friedman test for Gc vs G1 vs G2 vs G3 and a Wilcoxon signed-rank test for all four comparisons individually. To prove the hypothesis of RQ 1, we compare Gc vs G3 and for RQ 2 the other three comparisons.
@@ VI-A1 Learning Performance
The Friedman test shows significant differences among the groups in learning time($\chi^{2}(3)=18.725,p=0.003$). The average learning time shows that the participants spend the longest amount of time (all in minutes) when provided event-level causal demonstration in G1($M=24.71,SD=3.08$). To evaluate H1, we compare the Gc and G3.The average learning time of participants from G3 ($M=22.27,SD=3.23$) spend more time than those from Gc ($M=18.24,SD=2.60,p=0.001,Z=-2.981$). This shows that G3 which includes causal information increases the learning (training) time as compared to Gc which shows only the current task. Hence H1 is failed. For H4, we compare among G1, G2 and G3. The time spent in learning of G2 ($M=21.26,SD=2.71$) is not significantly different from that of G3 ($p=0.140,Z=-1.098$). Meanwhile, participants from G2 spend significantly less time in learning than those from G1 ($p=0.008,Z=-2.401$). Participants from G1 also spent significantly more time learning than those from G3 ($p=0.050,Z=-1.647$). We conclude that H4 is partly proved as the most time was taken by G1 and least by G2
@@ VI-A2 Testing Performance
We evaluate the remaining four hypotheses H2, H3, H5, and H6 in this section. The participant’s performance during the testing phases with respect to the average total task time is shown in  Figure 9 (H2 and H5). The number of mistakes of three different levels made during testing phase is demonstrated in  Figure 10 (H3 and H6).
The Friedman test shows significant differences among the groups in testing time ($\chi^{2}(3)=15.9,p=0.001$) The average testing time of the participants from G3($M=4.31,SD=0.35$) is significantly lower than all three other groups, namely Gc($M=5.20,SD=0.69,p=0.002,Z=-2.824$), G1 ($M=4.92,SD=0.47,p=0.001,Z=-2.981$), and G2 ($M=4.61,SD=0.24,p=0.036,Z=-1.804$). This proves H2 and H5. The average testing time of the participants from G1 is not significantly different from those from G2 ($p=0.058,Z=-1.5689$).
@@
We also evaluated the number of mistakes made by the participants during the testing phase. The Friedman tests show no significant difference among the groups in total mistake numbers($\chi^{2}(3)=3.3,p=0.348$), event-level mistake numbers ($\chi^{2}(3)=0.5,p=0.919$), interaction-level mistake numbers ($\chi^{2}(3)=5.825,p=0.120$), and gesture-level mistake numbers ($\chi^{2}(3)=4.525,p=0.210$). This shows that both H3 and H6 fails.
@@ VI-B1 Likert-Scale Questionnare
  • Understanding. The Friedman test shows a significant difference between all four groups for Q1($\chi^{2}(3)=8.325,p=0.040$) and Q2($\chi^{2}(3)=8.05,p=0.044$) but not Q3($\chi^{2}(3)=0.175,p=0.98$). We performed pairwise comparisons among the groups for Q1 and Q2. For Q1, G3 ($M=4.42,SD=0.75$) understands the importance of the task better as compared to Gc ($M=2.58,SD=1.24,p=0.005,Z=-2.548$). G3 ($M=3.92,SD=0.9$) also knows better the outcome of their actions beforehand as compared to Gc ($M=2.75,SD=1.35,p=0.038,Z=-1.765$)  Figure 11.
@@ VI-B2 NASA TLX
We used NASA TLX’s responses to evaluate the perceived subjective mental workload experience. The findings are present in  Figure 14. After confirming the non-normality with a Shapiro-Wilk test, a Friedman test was conducted on the results obtained from an unweighted NASA TLX to compare all the groups. None of the metrics showed significance: Mental Demand ($\chi^{2}(3)=3.875,p=0.275$), Physical Demand ($\chi^{2}(3)=1.8,p=0.614$), Temporal Demand($\chi^{2}(3)=4.225,p=0.238$), performance ($\chi^{2}(3)=2.625,p=0.453$), Effort($\chi^{2}(3)=3.525,p=0.317$) and frustrations($\chi^{2}(3)=2.125,p=0.546$). We further do not perform any adjustments in p-values for multiple comparisons, as there is no significant result as shown in  Figure 14.
@@ VI-C Research Question Analysis
  • RQ1: We observed that H1 and H3 did not yield successful outcomes, whereas H2 was confirmed to be accurate. H1 posited that the learning time in G3 would be less as compared to Gc. Yet, the learning time in G3 is significantly higher than that of Gc. After failing of two hypotheses we further analyzed more results. To explore performance improvements during the learning phase, we incorporated an additional metric based on the number of tutorial revisits [6]. Interestingly, the average revisits for Gc ($M=3.17,SD=0.72$) were significantly higher compared to G3 ($M=2.33,SD=0.49,p=0.018,Z=-2.090$), as depicted in  Figure 8. This shows the difficulty in learning the content. Regarding H3, it was inconclusive due to the absence of a significant difference. We conclude that the absence of significance in the number of mistakes is due to the similar mistakes made during testing, which led to close counts of mistakes across the groups at all three levels. The rejection of H1 and H3 along with the confirmation of H2 suggest the answer to RQ1 is that visualizing causality in task learning results in a shorter performance time in testing with the compromise of more time spent in learning while there is no evidence of fewer errors while performing the task.
@@
  • RQ2: We can see from the results that H4 is rejected although G1 did take the most time which corresponds to event-level causality. H5 was proven valid whereas H6 is not supported. From further analyzing the data we found that the reason behind the failure of H6 is the same as discussed for H3 in the above paragraph for the mistakes made in the task. To answer RQ2, we conclude from our study that among all three levels, visualizing only event-level causality introduces the most time spent in learning while visualizing causality at the interaction level costs the least time in learning. In addition, learning with gesture-level causality visualization results in the least completion time post-learning, while there is no evidence that learners would make fewer mistakes post-learning.
@@ VII-A Better Learning Results with Intention-driven Causality
VII-B Levels of Details: the Deeper, the Better
@@
From our cross-comparisons among G1, G2, and G3, we notice that both G2 and G3 have significantly less learning time than G1. The higher learning time of G1 comes from the fact that showing only event-level instruction makes it hard for the learners to memorize the procedure and results in the learners more frequently referring to the instructions. ”The instruction is too long for me to remember, and I have to rewind to the part I am doing (G1, P3).” The feedback from Memory - Q2 indicates that without learning the interaction and gesture/affordance levels of intention-driven causality, G1 has a poorer understanding of how to operate the tool. A similar conclusion can be drawn from Memory - Q3 where G1 gives lower ratings on their performance than G2 and G3 regarding their precision.
@@
Prior works [103, 8, 104] have shown that the more details provided to the users of the MR system, the higher cognitive load is forced on the users, resulting in decreasing learning performance. Yet, with details added in our study, we do not observe a significant increase in learning time or a significant difference in users’ subjective confidence about the task they have learned (Self-confidence - Q1,2,3). To sum up, we conclude that visualizing gesture-level causality information in manual task learning in MR provides the learners with the best understanding and memorization of the tasks and the highest post-learning time performance in the tasks. However, gesture-level causality visualization in manual task learning is not the best practice when efficiency in learning is the priority, because it results in more learning time compared to no-causality-visualization.
VII-C In-place 3D visualization
In our study, we employed a hybrid approach, combining 2D videos and 3D animations, to present content to learners about the tasks [3]. We utilized a hierarchical representation in our visualizations to align with the hierarchical mental models that humans naturally employ to decompose and understand tasks. Participants are prompted to look at 3D visualization for the current task, and then look at the hierarchical 2D visualization for complete causal information. The different placements of two visualizations can cause discontinuity in participants’ learning experiences.
A potentially more promising way to visualize both pieces of information and avoid discontinuity is via full 3D animations fixed at the objects showing current tasks followed by the causal tasks [105]. Additionally, other forms of visual cues [15, 24, 57] such as arrows, bounding boxes around the objects, and textual description can be integrated to provide a more designated learning purpose. Learners can better anticipate the implications of their actions within the 3D space [6] , leading to deeper learning